### v2

# Accepted to 2024 European Conference on Computer Vision

Abstract
Spiking Neural Networks (SNNs) offer a biologically inspired approach to computer vision that can lead to more efficient processing of visual data with reduced energy consumption. However, maintaining homeostasis within SNNs is challenging, as it requires continuous adjustment of neural responses to preserve equilibrium and optimal processing efficiency amidst diverse and often unpredictable input signals. In response to these challenges, we propose the Asynchronous Bioplausible Neuron (ABN), a dynamic spike firing mechanism that offers a simple yet potent auto-adjustment to variations in input signals. Its parameters, Membrane Gradient (MG), Threshold Retrospective Gradient (TRG), and Spike Efficiency (SE), make it stand out for its easy implementation, significant effectiveness, and proven reduction in power consumption, a key innovation demonstrated in our experiments. Comprehensive evaluation across various datasets demonstrates ABN’s enhanced performance in image classification and segmentation, maintenance of neural equilibrium, and energy efficiency.
I Introduction
Computer vision has experienced significant success over the last years due to the advancement of Artificial Neural Networks such as Convolutions Neural Networks (CNNs) and Vision Transformers. Recently, the attraction of Spiking Neural Networks (SNNs) has been growing, due to their potential in real-time sensory processing, closer imitation of biological brain functionalities, and low power consumption. The advent of event-based cameras further accelerates this evolution, offering a promising path to harness the strengths of SNNs and address the challenges faced by traditional vision systems. This combination paves the way for more energy-efficient and adaptive solutions in computer vision [1].
Overview of Standard Spiking Encoding and Membrane Threshold Methods
Our Encoding Free and Dynamic Membrane Threshold Method
Figure 1: Comparative Visualization of SNNs approaches: The top image presents the overview of the SNN methodology showcasing the conventional processes of asynchronous event capture, event-to-frame conversion, and fixed thresholding in neuronal spike response. In contrast, the bottom image illustrates an overview of the proposed method, highlighting the explicit encoding of events to spikes and implementing a dynamic thresholding mechanism for neuronal spiking. This juxtaposition underscores the novel methodology proposed.
SNNs’ performance benefits from homeostasis, i.e. maintaining a stable neuron activity level, neither spiking too often nor staying too quiet [2, 3]. To achieve this, standard methods adjust parameters like fixed firing thresholds or synaptic weights. These strategies require fine-tuning, and may not be versatile across network architectures. The fixed thresholds in spiking neurons, as shown in Fig. 1, can be sensitive to initial conditions, leading networks to be overly active or dormant. The inflexibility in adapting to varying inputs makes neurons either too sensitive or unresponsive, affecting negatively the performance. It also compromises energy efficiency, with either high power usage from excessive firing or missed data from sparse firing.
To address the above challenges, we introduce the Asynchronous Bioplausible Neuron (ABN) function, which can be incorporated into the two most prominent bioplausible neuron models: the Spike Response Model (SRM) [4] and the Leaky Integrate-and-Fire (LIF) model [5]. ABN is comprised of three novel elements MG, TRG, and SE. In the MG component, we are the first to propose calculating gradients to quantify the movement of membrane potential, akin to the ‘rate of membrane potential change’ observed in biological neurons, which reflects how neurons dynamically adjust their potential in response to inputs [6]. This provides the direction of the movement necessary to maintain the neuron’s output firing. In TRG, we are the first to implement inertia in threshold movement using a sliding window to control bursts of spikes, offering network-independent threshold control. This mirrors the ‘adaptive threshold’ or ‘activity-dependent threshold modulation’ observed in biological neurons, where firing thresholds adapt based on historical activity to maintain stability [7]. Although the principle of the SE element is adopted from existing literature, we are the first to implement it using a sliding window mechanism and asynchronous event data, relating it to ‘neuronal efficiency’ or ‘firing efficiency,’ which measures how effectively a neuron converts electrical inputs into action potentials, a crucial factor in preventing response saturation [8].
We validate the effectiveness of our approach using two SNNs for object segmentation and image classification tasks under both normal and various degraded conditions, such as low light, high occlusion, etc. Further, our method is tested on six different datasets: N-MNIST [9], ESD-1, ESD-2 [10], DVS128-Gesture [11], N-ImageNet [12] and CIFAR-10 DVS [13] to demonstrate its general applicability. Additionally, we conducted an extensive study of the homeostasis (i.e. stability test of neuron firing), an ablation study to understand the impact of elements of the proposed function, and compared the power consumption with state-of-the-art methods.
SOTA Thresholding MethodRefer to caption
Our MethodRefer to caption
Figure 2: Dynamic thresholding for SNNs: Existing methods are based on the LIF neuron which uses statistical cues for thresholding (a), however they result in unstable thresholds (b) and consequently low accuracy for segmentation methods (c). Our proposed method incorporates weighted inputs in the form of MG, TRG, and SE (d), which results in a smoothed dynamic threshold (e), and improved segmentation (f).
The remainder of this paper is structured as follows: Section II discusses related work, Section III details the proposed methodology, Section IV presents experimental results, and Section V concludes and discusses future research.
II Related Work
II-A Computer vision applications of SNNs
The exploration of SNNs for object segmentation and image classification has ignited widespread interest in the field, as the SNNs replicate the biological neuronal system to provide efficient and effective solutions for this task. Initial works in object segmentation leaned on feed-forward SNN architectures, utilizing the spike-timing-dependent plasticity (STDP) learning rule for unsupervised learning of features [14]. The Backpropagation Through Time (BPTT) method introduced in [15] for training SNNs, offers a more structured and supervised learning approach that significantly enhances feature extraction and classification accuracy, yet it is limited by its high computational complexity, substantial memory requirements, susceptibility to vanishing and exploding gradients, challenges in capturing long-term dependencies, and inefficiency in real-time learning applications. The introduction of Convolutional Spiking Neural Networks (CSNNs) marked a significant development in image classification. By incorporating convolutional layers into the network, these models could better handle the spatial complexities inherent in images, significantly enhancing segmentation performance. Concurrently, they maintained the temporal dynamics native to SNNs, enabling efficient processing of streaming data [16]. Recently, there has been a surge in the application of recurrent spiking neural networks (RSNNs) to image classification. These networks leverage feedback connections, enhancing the network’s ability to capture temporal dependencies in the data. Furthermore, their capacity to retain information over extended periods makes them particularly suited to segmentation tasks involving sequential or video data [17].
The memory-based methods such as those utilizing LSTM and RNN principles, highlighted by [18] and [19], offer advancements yet face limitations due to their application across entire networks, constraining individual neuron-based threshold modeling. Similarly, Spiking Neural Networks (SNNs) have been applied to tasks like segmentation and object detection with promising results as demonstrated by [20] and [21]; however, these models often draw from Artificial Neural Network (ANN) principles, potentially veering away from the bio-plausibility sought to closely mimic biological brain functions. These developments underscore the challenges in achieving nuanced modeling and bio-inspired design within neural network research, highlighting the need for approaches that balance technical efficacy with the intricacies of biological mechanisms. Generative networks [22] in event-based vision struggle due to the asynchronous and unpredictable nature of the data, leading to performance issues in scenarios with sporadic events. This challenge highlights the complexity of creating bio-plausible models that maintain stability and accuracy in the face of such data variability.
II-B Dynamic Thresholding in SNNs
Research on SNNs has delved deeply into dynamic thresholding methods to bolster adaptability. For instance, while predictive coding emphasizes anticipating input patterns, it overlooks the spike timing significance, an aspect STDP-based adjustments address [23], [24]. The holistic approach of network activity regulation contrasts with the inherent refractory period modulation, focusing on preventing rapid neuron firing [25]. Furthermore, spiking models like SRM and LIF have advanced neuron dynamics representations [4, 26]. However, gaps remain, especially in dynamic threshold modulation. [27] introduced a method utilizing a scaling factor to adjust the threshold in response to increased input. Yet, since it was tailored specifically for incremental spike membrane potential, it faltered when faced with variations in membrane decays. On the other hand, [28] adopted exponential functions to dictate decay. This approach, however, resulted in fluctuating thresholding levels, failing to ensure a smooth threshold curve. Meanwhile, [29] ventured into a firing counter-based threshold adjustment. While innovative, it necessitated meticulous fine-tuning and exhaustive experimentation, leaving the determination of an optimal count mired in ambiguity.
Previous research may not addressed all the dynamics in the input signal, especially in highly complex event-based vision signals. In the event-based vision, SNNs face unique challenges: Neurons within the same layer exhibit varying event counts; some neurons experience more events than others. Consequently, applying uniformly the BDETT method [30], which utilizes average membrane potential of the neurons in a layer to predict the threshold of all the neurons in that layer, fails to reflect the individual dynamics of each neuron. This paper addresses notable research gaps. No prior studies have demonstrated that a bioinspired dynamic threshold can stabilize spike firing in semantic segmentation tasks. Additionally, these studies typically employ simplistic datasets such as NMNIST [9] and rarely evaluate system robustness under challenging conditions like motion blur or low light, vital for assessing the system’s homeostasis [31].
III Methodology
III-A Pre-requisite
III-A1 Event-based cameras
Event-based vision cameras are designed to register variations in logarithmic light intensities by recording individual pixel-level changes, referred to as events. These events formulate a continuous asynchronous event stream, which can be mathematically articulated as a series of ordered tuples. A tuple corresponds to an event $i$ and includes its spatial coordinates $(x_{i},y_{i})$ in the image frame, timestamps $t_{i}$, and polarity value $z_{i}$ [32]:
${(x_{1},y_{1},t_{1},z_{1}),(x_{2},y_{2},t_{2},z_{2}),...,(x_{n},y_{n},t_{n},z_{n})}$ (1)
III-A2 Leaky integrate-and-fire
The LIF model, as shown in Fig. 1, is characterized by its linear integration of inputs until a threshold $\Theta(t)$ is reached, causing its firing [5]. Mathematically, the sub-threshold dynamics of the membrane potential $V(t)$ of an LIF neuron can be described by the following differential equation:
$\tau_{m}\frac{dV(t)}{dt}=-(V(t)-E_{L})+R_{m}I(t)$ (2)
where $\tau_{m}$ is the membrane time constant that dictates the speed at which the neuron responds to inputs, $E_{L}$ is the leak potential or resting potential representing the equilibrium potential of the neuron in the absence of any input, $R_{m}$ is the membrane resistance, and $I(t)$ represents the total synaptic input to the neuron that might come from other neurons or external sources. When the membrane potential $V(t)$ reaches the threshold $\Theta(t)$, the neuron emits a spike and then resets its potential to a reset value $V_{\text{reset}}$, typically below $\Theta(t)$. Additionally, after spiking, the neuron enters a refractory period for a duration $T_{\text{ref}}$ during which it cannot fire again. The "leaky" aspect of the LIF model comes from the term $(V(t)-E_{L})$ in the differential equation. This term ensures that in the absence of any input, the neuron’s potential will gradually revert to the leak potential $E_{L}$.
III-A3 Spike response model
In the SRM [4], incoming spike trains $s_{i}(t)$ are initially transformed into a spike response signal by convolving them with a spike response kernel $\varepsilon(\cdot)$, denoted as $(\varepsilon\ast s_{i})(t)$. Subsequently, this spike response signal is scaled by a synaptic weight $w_{i}$. The depolarization process involves summing up these scaled spike response signals across all incoming spikes, mathematically represented as $\Sigma_{i}w_{i}(\varepsilon\ast s_{i})(t)$. The SRM also incorporates a mechanism for hyperpolarization when a spike $s(t)$ is triggered, characterized by a refractory potential $(\zeta\ast s)(t)$, where $\zeta(\cdot)$ serves as the refractory kernel. This dual-component approach enables SRM to capture both excitatory and inhibitory dynamics in neuronal spike activity.
Figure 3: Schematic Diagram of ABN. The DVS camera outputs data in the form of asynchronous events. It is passed into a SNN with neurons in the first layer equal to the number of pixels on the camera. Each neuron adopts the proposed ABN where it takes the weighted input from the MG, TRG, and SE to control the dynamic threshold.
III-A4 Asynchronous events to spike conversion
Traditionally, rate coding and temporal coding [33], [34] are the popular methods to convert frames, from a traditional CMOS, into spiking trains. In this study, similar to [35], each event is treated as a spike to avoid the conversation of the input signal into a spike train. This approach essentially uses a form of temporal coding, because the information is conveyed through the timing of the spikes. We have chosen to drop the polarity aspect of the events due to its high sensitivity, which can introduce noise and complicate the signal processing. Mathematically, E(t) is the event stream, $e=(x,y,t)$ is an individual event with coordinates $x$, y and timestamp t, and $S_{i}(t)$ is the spike train for the i-th neuron:
$S_{i}(t)=\begin{cases}\delta(t-e_{t})&\text{if }e_{x}=x_{i}\text{ and }e_{y}=y_{i}\\ 0&\text{otherwise}\end{cases}$ (3)
Here, $\delta(t-e_{t})$ is the Dirac delta function, which represents a "spike" at time $t$.
III-B Asynchronous Bioplausible Neuron
The proposed ABN is a weighted combination of the novel Membrane Gradient (MG), Threshold Retrospective Gradient (TRG), and Spike Efficiency (SE) for threshold adjustment, uniquely tailored for each neuron. This individualization is crucial as each neuron receives a unique frequency of events, necessitating adaptation to input variations and ensuring the maintenance of firing stability, irrespective of the layer. Crucially, ABN’s design embodies simplicity and functions as a lightweight adjustment, seamlessly integrating into existing frameworks without the need for complex modifications. The details of each component are discussed in the subsequent subsections below. The mathematical equation of the ABN is:
$\Theta_{i}(t+1)=\Theta_{i}(t)+k_{1}MG_{i}-k_{2}TRG_{i}+k_{3}SE_{i}$ (4)
The equation models the dynamic adjustment of the neuron’s firing threshold at the next time step $t+1$ based on various factors. In this equation, $\Theta(t)$ is the existing threshold at time $t$, and $k_{1},k_{2},k_{3}$ are weights that determine the relative influence of each term. The term $k_{1}\cdot MG_{i}$ increases the threshold when the membrane potential is rapidly changing, preventing the neuron from firing too easily. Conversely, the term $-k_{2}\cdot TRG_{i}$ decreases the threshold if there has been a historical trend of rapid threshold increases. Lastly, the term $k_{3}\cdot SE_{i}$ takes into account the neuron’s recent spiking activity. Specifically, if the neuron is overly active, this term will make it more difficult for the neuron to fire in the future, thereby ensuring efficiency. Overall, this equation encapsulates a multi-faceted approach to dynamically adapting the neuron’s firing threshold.
III-B1 Membrane Gradient (MG)
The rate of change of the membrane potential is a critical parameter for predicting a dynamic threshold in SNNs. It serves as an indicator of the neuron’s responsiveness to incoming spikes and plays an essential role in governing the neuron’s firing behavior. Positive correlations between dynamic thresholds and membrane potentials have been observed in several areas of diverse biological nervous systems. In [36], the average membrane potential is correlated with an average threshold to predict the dynamic threshold. Instead, in this work we correlate the rate of change of membrane potential with a threshold. A high rate of change often signifies that the neuron is receiving a burst of spikes, necessitating a higher threshold to maintain neuronal homeostasis. Conversely, a lower rate of change could indicate that the neuron is less active, allowing for a lower dynamic threshold to be sensitive to sparse inputs. Incorporating the rate of change of the membrane potential into the calculation of a dynamic threshold ensures that the SNNs can adapt to varying synaptic activities, thereby emulating the adaptability and energy efficiency observed in biological neurons.
The BDETT method presented in [30] calculates the dynamic threshold using the average of membrane potentials and thresholds. In a layer with varied neuronal activity, not all neurons are always active. Including inactive neurons in this calculation can pull the average value lower. Applying this reduced average threshold to highly active neurons may result in excessive spiking, leading to overly aggressive neuronal activity. In contrast, in this work, we adopt the neuron-specific derivative term, denoted as $\eta\cdot\frac{dv_{i}(t)}{dt}$, which captures the rate of change of membrane potential of neuron $i$ at time $t$, denoted by $v_{i}(t)$. To numerically compute the derivative, a finite difference approximation is employed. This involves the subtraction of the membrane potential at the previous timestep from the current potential, followed by division by the timestep size $\Delta t$. The scaling factor $\eta$, which is set to a value less than 1, reduces the rate of change to ensure that the rate of change of the threshold remains less than the rate of change of the potential. The explicit equation for the derivative term is given as:
$MG_{i}=\frac{dv_{i}(t)}{dt}\approx\frac{v_{i}(t)-v_{i}(t-1)}{\Delta t}$ (5)
III-B2 Threshold Retrospective Gradient (TRG)
Biological systems exhibit a negative correlation with the preceding rate of depolarization, which indicates the excitatory status. We propose considering a historical rate of change for the threshold $\Theta(t)$. This is because focusing solely on the preceding rate of depolarization as implemented in [30] is insufficient and can result in an unstable or overly sensitive dynamic threshold. This approach lacks memory, reacting too quickly to transient changes without considering longer-term neuronal activity patterns. Consequently, this can result in excessive threshold fluctuations and potentially amplify noise, undermining network stability and reliability. Therefore, we take into account the average of the historical rate of change of the dynamic threshold. Utilizing the historical rate of change of the dynamic threshold to calculate the current threshold can provide the neuron with inertia or memory of its past states. This can be useful in situations where the neuron should react not only based on the current state but also considering its past behavior. In discrete terms, if $\Theta(t)$ is the threshold at time $t$, then the rate of change at $t$ is:
$G_{i}(t)=\frac{\Theta_{i}(t)-\Theta_{i}(t-1)}{\Delta t}$ (6)
The historical rate of change is an average of these rates over a recent window of time, $N$ time steps:
$\overline{G}_{i}(t)=\frac{1}{N}\sum_{j=0}^{N-1}G_{i}(t-j)$ (7)
In order to ensure that the older rates influence the current threshold less within the sliding window, a decay factor $\alpha,(0<\alpha<1)$, is used:
$TRG_{i}=\frac{1}{N}\sum_{j=0}^{N-1}\alpha^{j}G_{i}(t-j)$ (8)
with values closer to 0 suppressing more the older rates.
In summary, leveraging the historical rate of change of the dynamic threshold has introduced a form of temporal inertia to the neuron, creating a relationship to patterns in input data over time. Importantly, it is especially useful in scenarios where temporal patterns are crucial.
III-B3 Spike Efficiency (SE)
The ratio of spike output $C_{\text{out}}$ to spike input $C_{\text{in}}$ can serve as a crucial metric for determining the dynamic threshold in SNNs. Intuitively, this ratio indicates how efficiently a neuron converts incoming spikes into outgoing spikes. A high SE suggests that the neuron is highly responsive to incoming activity, and could therefore benefit from a high dynamic threshold to maintain energy efficiency and reduce the likelihood of over-saturation. Conversely, a low SE implies that the neuron is less responsive, possibly requiring to lower a dynamic threshold to avoid unnecessary firing and thereby preserve neuronal and computational resources. This ratio is mathematically represented as:
$r_{i}(t)=\frac{C_{\text{out},i}(t)}{C_{\text{in},i}(t)}$ (9)
When using a sliding window method, we only consider $N$ previous time steps to calculate $C_{\text{in}}$ and $C_{\text{out}}$. This allows the threshold to adapt quickly to recent changes in spike rates.
$SE_{i}=\bar{r}_{i}(t)=\frac{\sum_{j=t-N}^{t}C_{\text{out},i}(j)}{\sum_{j=t-N}^{t}C_{\text{in},i}(j)}$ (10)
Incorporating SE into the dynamic threshold calculation enables the SNN to adjust its sensitivity according to the context of incoming activity. This inclusion ensures a more nuanced and adaptive network behaviuor, aligning closely with biological plausibility while maintaining computational efficiency. It allows the network to operate in a balanced regime, irrespective of the incoming spike rates, ensuring robust and stable performance across different input conditions. This adaptive mechanism is invaluable, especially when individual neurons within a network necessitate autonomous regulation of their activity, contingent on their functional significance or the unique input patterns they discern.
IV Results
IV-A Experimental Protocol
We assess the effectiveness of the proposed ABN in two Computer Vision tasks: Image Classification for the N-MNIST [9], CIFAR10-DVS [13], N-ImageNet [12], DVS128 gesture recognition [11] datasets, and Object Semantic Segmentation for the ESD-1 [10] and ESD-2 [10] datasets. The details of the datasets are explained in supplementary section B. The Spiking MLP employs Spatio-Temporal Backpropagation [30] for training, allowing the network to consider both spatial and temporal information during the learning process. Although the input to the first layer should be a spike train, we can directly use asynchronous event-based vision data without requiring specific conversion. The initialization of network parameters, such as weights and thresholds, is crucial for stabilizing the network’s firing activities. We adopt the initialization strategy as discussed in [37]. To make a fair comparison with [30], we have used a similar experimental configuration. The model was trained on a GPU PC featuring 128GB of memory, an Intel Xeon W-2155 Processor, and NVIDIA Quadro RTX 8000 48GB graphics cards, housed in a Lenovo ThinkStation P520. While this is essential for assessing the efficacy of the proposed method, implementing it directly on neuromorphic hardware may require additional engineering efforts, which are beyond the scope of this work. Custom code was written to develop the LIF and SRM within this framework. Synaptic weights are initialized randomly, and a decay constant, $\tau_{1}$, is set to 0.1 ms. The weights for constants $k_{1}$, $k_{2}$, and $k_{3}$ are experimentally selected as 0.25, 0.50, and 0.25, respectively (see supplementary document).
IV-B Comparative Results across Diverse Datasets
TABLE I: Classification Accuracy across Various Datasets and Models
Method N-MNIST DVS128 CIFAR10 SRM LIF SRM LIF SRM LIF Spiking RBM [38] 92.1 93.16 87.38 90.21 83.04 86.25 Spiking MLP (BP) [39] 94.52 97.66 90.72 93.4 86.23 90.06 Spiking MLP (STDP) [40] 93.47 95 90.8 92.01 87.5 91.7 Spiking MLP (STBP) [37] 97.13 98.89 92.54 93.64 86.23 91.73 DT1 [27] 99.05 99.4 95.01 96.88 89.17 92.65 DT2 [29] 98.13 98.24 92.54 95.54 89.38 91.47 BDETT [30] 99.15 99.45 94.09 96.05 91.61 93.5 ABN (Ours) 99.23 99.48 95.64 98.74 93.5 94.74
TABLE II: Segmentation Accuracy across Various Datasets and Models
Method ESD-1 ESD-2 SRM LIF SRM LIF Spiking RBM [38] 48.95 51.05 45.2 50.31 Spiking MLP (BP) [39] 49.53 54 45.5 52.42 Spiking MLP (STDP) [40] 52.63 58.84 46.7 49.44 Spiking MLP (STBP) [37] 55.82 63.09 49.07 51.37 DT1 [27] 57.55 61.45 50.13 54.59 DT2 [29] 58.28 64.3 53.49 56.5 BDETT [30] 61.02 65.39 51.36 55.46 ABN (Ours) 61.56 67.29 56.23 58.04
The evaluation of our ABN method’s generalization capabilities, across diverse datasets and neuron models is shown in Table I and II. Performance on the classification datasets was evaluated using classification accuracy and on the object segmentation using event accuracy. The metrics are detailed in the supplementary document section C. On the N-MNIST dataset, ABN achieved the highest classification accuracy (99.23% with SRM and 99.48% with LIF), surpassing other methods like BDETT and DT1. In the ESD-1 and ESD-2 datasets, designed for robotic grasping tasks, ABN again outperformed others, showing particularly strong performance in the ESD-2 dataset with unknown objects (56.23% with SRM and 58.04% with LIF). For the DVS128 Gesture, CIFAR10-DVS and N-ImageNet datasets, ABN recorded the highest scores 95.64%, 93.5% and 56.73% with SRM and in LIF 98.74%, 94.74% and 57.04% , indicating its robustness and effectiveness. This comprehensive performance underscores the superior adaptability and effectiveness of the ABN method across a variety of datasets and computer vision tasks.
IV-C Evaluation on Degraded Input on ESD-1
TABLE III: Segmentation accuracy of known objects in various conditions.
Exp 1: varying clutter objects, Bright light, 62cm height, Rotational motion, 0.15 m/s speed Method 2 Obj 4 Obj 6 Obj 8 Obj 10 Obj Spiking MLP (STBP) [37] 74.16% 73.05% 66.79% 66.42% 54.42% DT1 [27] 77.56% 74.65% 68.20% 68.70% 55.49% DT2 [29] 78.46% 75.42% 69.28% 69.62% 56.16% BDETT [30] 79.07% 78.03% 74.14% 69.83% 58.28% ABN (ours) 81.49% 79.74% 74.97% 72.16% 60.03% Exp 2: 6 Objects, varying lighting conditions, 62cm height, Rotational Motion, 0.15 m/s speed. Method Bright Light Low light Spiking MLP (STBP) [37] 58.73% 61.59% DT1 [27] 61.32% 62.91% DT2 [29] 61.24% 64.53% BDETT [30] 63.92% 66.03% ABN (Ours) 65.98% 67.42% Exp 3: 6 Objects, Bright Light, 62cm height, Varying directions of motion, 0.15 m/s speed. Method Linear Rotational Partial Rotational Spiking MLP (STBP) [37] 48.09% 63.99% 67.43% DT1 [27] 49.09% 65.14% 69.31% DT2 [29] 50.43% 66.89% 69.79% BDETT [30] 52.09% 68.60% 72.04% ABN (Ours) 55.50% 69.45% 73.54% Exp 4: 6 Objects, Bright Light, 62cm height, Rotational motion, Varying speed. Method 0.15 m/s 0.3 m/s 0.1 m/s Spiking MLP (STBP) [37] 53.16% 68.12% 56.29% DT1 [27] 54.56% 70.27% 60.16% DT2 [29] 55.27% 72.31% 59.63% BDETT [30] 56.67% 75.09% 63.16% ABN (Ours) 59.63% 76.72% 63.56% Exp 5: 6 Objects, Bright Light, Varying camera height, Rotational motion, Varying speed. Method 62 cm 82 cm Spiking MLP (STBP) [37] 62.93% 55.05% DT1 [27] 63.82% 59.08% DT2 [29] 65.63% 59.93% BDETT [30] 65.82% 59.00% ABN (ours) 68.40% 63.27%
We conducted an experiment focused on evaluating the segmentation performance of various state-of-the-art methods, and compared them to our proposed ABN, under conditions of degraded inputs. This experiment was essential to mirror real-world scenarios affected by factors like low light, occlusion, variations in speed and distance from the object, and the direction of camera motion. These factors significantly impact the quality of input, making it crucial to determine the robustness of the proposed method in less-than-ideal conditions. Utilizing the ESD-1 dataset, we measured the event accuracy to gauge the performance. The results across various conditions, as depicted in Table III, consistently show that the ABN method outperforms all other SNN methods in all tested scenarios. Clearly, the superior ability of ABN to maintain high segmentation accuracy despite the challenging input conditions, confirms its efficacy and potential for real-world applications.
IV-D Homeostasis Evaluation
Homeostasis in SNNs is crucial as it reflects the network’s ability to maintain stable internal conditions despite external environmental variations. It is an important indicator of the network’s reliability and efficiency, especially when dynamic input is considered.
TABLE IV: Homeostasis Performance on ESD-1 Known Object
$FR\_m$ $\Delta FR\_m$ $FR\_m\_std$ $\Delta FR\_m\_std$ $FR\_s\_std$ $\Delta FR\_s\_std$ Ideal Condition Spiking MLP (STBP) [37] 0.631 0.485 0.001564 DT1 [27] 0.524 0.459 0.001146 DT2 [29] 0.507 0.416 0.000983 BDETT [30] 0.436 0.403 0.001224 ABN (Ours) 0.385 0.253 0.000891 Occlusion Spiking MLP (STBP) [37] 0.783 0.152 0.582 0.097 0.001396 0.000168 DT1 [27] 0.692 0.168 0.561 0.102 0.003963 0.002817 DT2 [29] 0.586 0.079 0.484 0.068 0.002724 0.001741 BDETT [30] 0.47 0.034 0.357 0.046 0.002291 0.001067 ABN (Ours) 0.373 0.012 0.291 0.038 0.000784 0.000107 Light Spiking MLP (STBP)[37] 0.786 0.155 0.594 0.109 0.003942 0.002378 DT1 [27] 0.698 0.174 0.405 0.054 0.002508 0.001362 DT2 [29] 0.594 0.087 0.489 0.073 0.003641 0.002658 BDETT [30] 0.498 0.062 0.467 0.064 0.001157 0.000067 ABN (Ours) 0.405 0.02 0.295 0.042 0.000889 0.000002 Speed Spiking MLP (STBP) [37] 0.746 0.115 0.582 0.097 0.001666 0.000102 DT1 [27] 0.62 0.096 0.406 0.053 0.001167 0.000021 DT2 [29] 0.609 0.102 0.351 0.065 0.002839 0.001856 BDETT [30] 0.517 0.081 0.352 0.051 0.002077 0.000853 ABN (Ours) 0.305 0.08 0.206 0.047 0.000881 0.00001 Motion Direction Spiking MLP (STBP) [37] 0.701 0.07 0.505 0.02 0.003798 0.002234 DT1 [27] 0.601 0.077 0.409 0.05 0.003461 0.002315 DT2 [29] 0.415 0.092 0.397 0.019 0.002827 0.001844 BDETT[30] 0.395 0.041 0.351 0.052 0.00133 0.000106 ABN (Ours) 0.378 0.007 0.243 0.01 0.000875 0.000016 Size Variance Spiking MLP (STBP) [37] 0.759 0.128 0.533 0.048 0.003901 0.002337 DT1 [27] 0.675 0.151 0.449 0.01 0.002263 0.001117 DT2 [29] 0.551 0.044 0.45 0.034 0.001346 0.000363 BDETT [30] 0.481 0.045 0.354 0.049 0.002763 0.001539 ABN (Ours) 0.373 0.012 0.255 0.002 0.000879 0.000012
In an experimental setup across all variations of the ESD-1 dataset, various statistical indicators were used to quantify the homeostasis of the host SNNs. Specifically, for a fair comparison with [30] we employed $FR_{m}$, $FR_{m\_std}$, and $FR_{s\_std}$ metrics, representing the average neuron firing rate over all trials, the mean of standard deviations of neuron firing rates over individual trials, and the variability of these standard deviations, respectively. Our proposed ABN model demonstrates significantly lower values across all three metrics and all conditions. These results suggest a stable and well-regulated network, highlighting the ABN model’s capacity for homeostasis.
IV-E Energy Consumption
Table V presents an analysis of power consumption, comparing CNN, SotA SNNs and the ABN using the DVS128 gesture recognition dataset. As proposed in [31], MAC (Multiply-Accumulate operations) and AC (Accumulate Count) were used to estimate power consumption. A detailed description of power consumption calculations is provided in the supplementary document.
CNN models such as [41] and [42], show a higher power consumption of 0.541W and 1.56W, respectively, despite their zero AC. In contrast, using SNN models leads to significantly lower power usage, with hybrid and SNN models [31] consuming 0.404W and 0.053W, respectively. BDETT’s SNN model [31] further reduces this consumption to 0.046W. Notably, our proposed SNN model outperforms these by achieving the lowest power consumption of just 0.038W, even with a higher AC of 105, underscoring the efficiency of our model in terms of energy usage compared to both CNNs and other SNNs.
TABLE V: Power Consumption Comparison
Method Model MAC AC Power (W) Calabrese [41] CNN 285 0 0.541 Baldwin [42] CNN 963 0 1.56 Asude [31] Hybrid 235 81 0.404 Asude [31] SNN 0.6 123 0.053 BDETT [30] SNN 0.7 115 0.046 Ours SNN 0.4 105 0.038
V Conclusion
This paper introduces the Asynchronous Bioplausible Neuron (ABN) within a spiking multi-layer perceptron (MLP). This innovative spiking neuron incorporates a dynamic threshold based on Membrane Gradient (MG) for Spike Frequency Adaptation, Threshold Retrospective Gradient (TRG) for Burst Suppression, Spike Efficiency (SE) for Homeostasis. The spiking MLP, which utilizes STDP, serves as the architecture for implementing the proposed neuron.
A methodology for object segmentation and image classification that employs a dynamic vision sensor and the ABN is also proposed. Remarkably, this system can directly process asynchronous, event-based vision signals without requiring any preprocessing. Its performance was rigorously evaluated on both conventional datasets (N-MNIST, DVS128 Gesture, N-ImageNet and CIFAR10-DVS) and specialized robotic grasping datasets (ESD-1, ESD-2). The method achieved state-of-the-art performance in terms of classification accuracy and event-wise accuracy across different degraded signal conditions. Additionally, our approach has demonstrated excellent homeostatic properties, maintaining stable neuronal firing rates under challenging conditions such as occlusions, low light levels, the presence of small or rapidly moving objects, and linear motion. Notably, it also exhibited the lowest power consumption among the models evaluated.
Our model effectively emulates certain biological neuron behaviours, but it does not fully capture all human neuron characteristics, highlighting an area for enhancement to improve biological authenticity. Future work could explore advanced dynamic threshold functions incorporating elements like dynamic refractory periods and network activity. Additionally, current neuromorphic hardware architectures are primarily restricted to Leaky Integrate-and-Fire (LIF) and Spike Response Model (SRM) configurations, which precludes the direct deployment of our algorithm on hardware directly. Addressing this limitation remains a pivotal area for future research to ensure robust real-world applicability.
Acknowledgements
This work was supported by Kingston University, the Advanced Research and Innovation Center (ARIC), which is jointly funded by Khalifa University of Science and Technology, STRATA Manufacturing PJSC (a Mubadala company) and Sandooq Al Watan under Grant SWARD-S22-015. In addition, the authors would like to extend their appreciation to Ipsotek, an Eviden Company, for their generous support and valuable contributions to the research project.